the length of a rectangle is 3 less the the width. the perimeter is 68 . what is the length and width of the rectangle ?
how do i do that equation?
how do i set up the equation?

Answer by checkley77(12844)

L=W-3
P=2L+2W
68=2(W-3)+2W
68=2W-6+2W
68=4W-6
4W=68+6
4W=74
W=74/4
W=18.5 ANS. FOR THE WIDTH.
L=18.5-3=15.5 ANS. FOR THE LENGTH.
PROOF:
68=2*18.5+2*15.5
68=37+31
68=68